of two wheels, the circumference of the large one is 3.5 times that of the smaller. In rolling a distance of 63 meters the smaller wheel makes ten more turns than the larger wheel. What is the circumference of each wheel?
ps can you please get back to me as quick as possible

Answer by rothauserc(4718)

let x be the circumference of the smaller wheel and 3.5x is the circumference of larger wheel
let n be the number of revolutions for the larger wheel
63 = n3.5x
63 = (n+10)x
solve for x in second equation and substitute in first
x = 63/(n+10)
63 = n(3.5 * 63/(n+10) )
63 = n (220.5/(n+10))
cross multiply fractions
63n+630 = 220.5n
630 = 157.5n
n = 4
substitute for n in second equation
63 = (4+10)x
x = 63/14 = 4.5 meters
circumference of smaller wheel is 4.5 meters and
circumference of larger wheel is 4.5 * 3.5 = 15.75 meters